# Semantic Segmentation Algorithm Based on Light Field and LiDAR Fusion

### v1

@@ III-B1 Point-Pixel Feature Fusion Module (PFFM)
Considering the characteristics of point cloud projections on the image, where the point cloud is densely concentrated in the central region of the image plane, and there are no point clouds in the upper and lower areas, we design the following interpolation method. First, we find the minimum bounding rectangle $M$ for all the projected point clouds on the image plane and let the feature points that can be projected within this rectangle be denoted as $P_{in}^{point}\in\mathbb{R}^{N\times c_{p}}$. There are no matching point clouds for pixels outside the rectangle. At this point, the point cloud feature map projected onto the image plane is given by $F_{point}^{img}=\left[P_{in}^{point},0\right]$. For the pixel positions within the rectangle that do not have corresponding point cloud projections, we perform feature interpolation as follows.
Let the pixel coordinates of the valid projected points $P_{in}^{point}$ be: $(x_{min},y_{min})=\min(x_{indices}),\min(y_{indices})$ and $(x_{max},y_{max})=\max(x_{indices}),\max(y_{indices})$. We define the range of the minimum bounding rectangle $M$ as $x_{min},x_{max}]\times[y_{min},y_{max}]$.
@@
$\text{mask}(x,y)=\begin{cases}1&(x,y)\in\{(x_{indices},y_{indices})\}\\ 0&(x,y)\notin\{(x_{indices},y_{indices})\}\end{cases}^{\top},$ (3)
@@
For each unassigned point $(x,y)$, we find its three nearest valid points $\{(x_{i},y_{i})\}_{i=1}^{3}$, calculate the interpolation weights, where the weights are inversely proportional to the distance between the unassigned point and The valid points are: $w_{i}=\frac{1}{d_{i}+\epsilon}$, where $d_{i}=\sqrt{(x-x_{i})^{2}+(y-y_{i})^{2}}$. We normalize the weights as $\hat{w}_{i}=\frac{w_{i}}{\sum_{j=1}^{3}w_{j}}$. The interpolated features are then:
$\hat{F}_{point}^{img}=\sum_{i=1}^{3}\hat{w}_{i}\cdot F_{point}^{img}(x_{i},y_{i})\quad,$ (4)
@@
$\hat{F}^{img}_{point}(x,y)=\begin{cases}{\scriptstyle\sum_{i=1}^{3}\frac{d_{i}+\epsilon}{\sum_{j=1}^{3}d_{j}+\epsilon}\cdot F^{img}_{point}(x_{i},y_{i}),\quad\text{mask}(x,y)=1}\\[5.0pt] F^{img}_{point}(x,y),\quad\text{mask}(x,y)=0\end{cases}^{\top},$ (5)
This gives us the point cloud projection in the feature map of the image plane within the bounding rectangle $M$. For regions outside the rectangle, we fill them with the corresponding image feature $F_{img}$ to obtain the complete point cloud projection feature map filled point $F_{fill\_point}$ is given by $F_{img}(x,y)$ if $(x,y)\notin M$, and $\hat{F}_{point}^{img}(x,y)$ if $(x,y)\in M$. Since the point cloud and image data features are in different spaces and have different network structures, they are not in the same feature space. To facilitate alignment of the point cloud projection feature map with the image feature space for full fusion, we design an alignment loss function(Pixel2point Loss). This loss function minimizes the difference between the feature spaces of the point cloud and image, enabling better fusion of the two. The alignment loss uses Mean Squared Error (MSE) to measure the difference between the point cloud feature map and the image feature map:
$\mathcal{L}_{align}=\frac{1}{N}\sum_{(x,y)}\left\|F_{fill\_point}(x,y)-F_{img}(x,y)\right\|_{2}^{2}\quad,$ (6)
After obtaining the complete point cloud projection feature map $F_{fill\_point}$, we concatenate it with the image feature map $F_{img}$ to create the initial fused feature map:
$F_{fused}=concat\left([F_{fill\_point},F_{\text{img}}]\right)\quad,$ (7)
Here, $concat(\cdot)$ denotes the concatenation operation.
We then apply a self-attention mechanism to the initial fused feature map to further refine the fusion, resulting in the final fused feature map $\hat{F}_{fused}$. The goal of the self-attention mechanism is fourfold: 1. To enhance the attention of the image feature map to itself. 2. To enhance the attention of the point cloud projection feature map to itself. 3. To allow the image feature map to gather useful information from the point cloud projection feature map. 4. To allow the point cloud projection feature map to gather useful information from the image feature map. We generate the Query, Key, and Value matrices from the fused feature map $\hat{F}_{fused}$: $Q=F_{fused}\cdot W_{Q}\in\mathbb{R}^{C_{q}\times(h\times w)}$, $K=F_{fused}\cdot W_{K}\in\mathbb{R}^{C_{k}\times(h\times w)}$, and $V=F_{fused}\cdot W_{V}\in\mathbb{R}^{C_{v}\times(h\times w)}$, where $W_{Q}\in\mathbb{R}^{C\times C_{q}}$, $W_{K}\in\mathbb{R}^{C\times C_{k}}$, and $W_{V}\in\mathbb{R}^{C\times C_{v}}$ are learnable projection matrices. The attention weights are calculated based on the similarity between the Query and Key:
$Attention(Q,K,V)=Softmax\left(\frac{Q^{T}K}{\sqrt{C_{k}}}\right)V^{T}\in\mathbb{R}^{(H\times W)\times C_{v}},$ (8)
@@
$\hat{F}_{fused}=\text{LayerNorm}\left(\text{Reshape}\left(Attention(Q,K,V)\right)\right),$ (9)
@@
$L_{fused\_img}=CE(\hat{y}_{fused},y_{gt})\quad,$ (10)
@@ III-B2 Depth Difference Perception Module (DDPM)
For occluded objects, fusing point cloud and image data may lead to conflicting features, as a single perspective alone cannot effectively supplement the information of the occluded objects. Thus, it is essential to assist the network in identifying these ’conflicting’ features. We hypothesize that when an object is occluded in the image but not in the point cloud, the depth of the object in the image should differ from its depth in the point cloud. Similarly, if an object is visible in the image but occluded in the point cloud, the depth of the object in the image should also differ from that in the point cloud. Therefore, we propose leveraging depth priors to guide the network in detecting regions with depth discrepancies.
For the input light filed images $\{\mathcal{L}_{1},\mathcal{L}_{2},\cdots,\mathcal{L}_{n}\}$, we use Zoe[46], a state-of-the-art deep learning model that leverages convolutional neural networks to predict depth maps from monocular images by learning spatial features and contextual information, to estimate the absolute depth of the images. Through Zoe, we obtain the predicted depth map for each image, denoted as $D_{pred}$. Meanwhile, the LiDAR can acquire 3D spatial coordinates for each point in the point cloud, as described in 2. The 3D coordinates of each LiDAR point can be projected onto the image plane to obtain the corresponding pixel coordinates $(u_{i},v_{i})$ and depth value $z_{i}$. These depth values are then filled into the corresponding positions on the image plane, generating the sparse depth map $D_{sparse}$. For each projected At this point, we have $D_{sparse}(u_{i},v_{i})=z_{i}$ for locations with projection points, and $D_{sparse}(u_{i},v_{i})=0$ for locations without projection points.
We only consider the positions with projection points. The difference map $D_{\text{diff}}$ is obtained by comparing the real sparse depth map and the predicted depth map, serving as the criterion for identifying occluded regions.
Considering that the accuracy of depth prediction from images dramatically decreases as the actual depth increases, we apply logarithmic smoothing to exaggerate the difference at close distances while minimizing the difference at farther distances. That is:
$D_{diff}(i,j)$ $=\log(D_{pred}(i,j)+\epsilon_{i,j})-\log(D_{gt}(i,j)+\epsilon_{i,j})$
$\quad\text{for}\quad(i,j)\in\mathcal{V}_{0},\quad\text{for}\quad(i,j)\notin\mathcal{V}$ (11)
where $\mathcal{V}$ represents the region with LiDAR projection points and $\epsilon=1\times 10^{-8}$. To map the depth differences to the same feature space as the network, we apply a two-layer convolutional neural network to $D_{diff}$:
$\hat{D}_{diff}=Conv(Conv(D_{diff}))\quad,$ (12)
The resulting difference feature map $\hat{D}_{diff}$ is then added to the attention map from Point-Pixel Feature Fusion Module, with the goal of guiding the network to focus on the occluded areas as much as possible, resulting in the final attention map:
$\text{Attention}_{final}=\hat{D}_{diff}+Attention(Q,K,V)\quad,$ (13)
This is then incorporated into the network to enhance the perception of occluded objects.
For the image branch, we use the fused feature loss $L_{fused\_img}$ and the image features $F_{img}$, which are upsampled to obtain the fused output $\hat{y}_{img}$. The loss is calculated as:
$L_{img}=CE(\hat{y}_{img},y_{gt})\quad,$ (14)
For a single image input, the total loss is:
$L_{img\_total}=L_{img}+L_{img\_lovasz}+L_{fused\_img}+\mathcal{L}_{align},$ (15)
For multiple image inputs, the total loss becomes:
$L_{img\_total}=L_{img\_center}+L_{img\_lvcenter}+L_{img\_lovasz}+\mathcal{L}_{align}\\ +L_{fused\_img}+\alpha_{1}\cdot\sum_{i=1}^{n}L_{img\_i}+\alpha_{2}\cdot\sum_{i=1}^{n}L_{img\_lvi},$ (16)
For the point cloud branch, we combine point-level features with image features and apply the segmentation head for point cloud predictions. The voxel features $F_{\text{voxel}}^{l}$ are also upsampled for segmentation. The point cloud and voxel losses are calculated as:
$L_{point}=CE(\hat{y}_{point},y_{point})+L_{lovasz}(\hat{y}_{point},y_{point}),$ (17)
$L_{voxel}=CE(\hat{y}_{voxel},y_{voxel})+L_{lovasz}(\hat{y}_{voxel},y_{voxel}),$ (18)
The total loss for the point cloud is:
$L_{point\_total}=L_{point}+L_{voxel}\quad,$ (19)
The overall network loss is:
$L_{total}=L_{img\_total}+L_{point\_total}\quad,$ (20)
@@ IV Experimental Results
To validate the effectiveness of our proposed the multimodal fusion-based segmentation approach, we conduct experiments on the TrafficScene dataset. All experiments were conducted on a server with an Intel 6330 CPU, 1.0 TB memory, Ubuntu 22.04.5 and CUDA version 12.2. The dataset is split into training, validation, and test sets at a 7:1:2 ratio (3924/594/1116 light filed images, 436/66/124 point clouds), using stratified sampling to balance category distributions.
IV-A1 Experimental Results
To validate the effectiveness of our dataset and multimodal fusion method, we conducted extensive experiments. First, we assessed the dataset’s validity by applying established image and light-field semantic segmentation methods, evaluating performance using mean intersection over union (mIoU).
@@ IV-A1 Experimental Results
CMNeXt [28] $\times$ $\times$ $\checkmark$ 83.26
@@
Mlpfseg (light filed images) $\times$ $\checkmark$ $\checkmark$ 84.97(+3.75) 92.38(+2.38)
Figure 5: mIoU for PSPNet, PSPNet LGA and Mlpfseg on small objects across all viewing angle
(a) Mlpfseg can correctly identify occluded bike
(b) Mlpfseg can correctly identify occluded pedestrian
Figure 6: Visualization of the results of different algorithms for occluded objects
@@
For point cloud segmentation, we benchmarked two pure point cloud methods, SPVCNN[24] and MinkowskiNet[23], alongside three multimodal fusion-based methods: PMF[41], 2DPASS[26], and Mseg3D[27]. Additionally, we evaluated our proposed Mipfseg and Mlpfseg. The final results are presented in TABLE I.
@@
The results show that attention-based methods outperform convolution-based ones in single-image segmentation, as attention mechanisms adaptively focus on crucial regions. CMNeXt[28] achieves the highest single-image mIoU of 83.15, while incorporating additional sub-aperture images in light-field-based methods further improves segmentation, surpassing the state-of-the-art (SOTA) with a mIoU of 83.61. This highlights the contribution of multiple viewpoints to central-perspective segmentation quality. For point cloud segmentation, Mseg3D[27] outperforms other methods, demonstrating the benefits of image-assisted segmentation. However, projection-based methods like PMF[41] suffer from 3D structure loss, resulting in inferior performance compared to pure point cloud methods such as SPVCNN[24] and MinkowskiNet[23].
@@
Our proposed Mlpfseg surpass existing SOTA methods in both image and point cloud segmentation. By interpolating image features, they alleviate the sparsity issues caused by point cloud projection. The attention mechanism enhances feature extraction for the image branch, while the depth-difference perception module improves occluded object segmentation, leading to overall performance gains in both modalities.
For both image and point cloud branches, better feature extraction from the image branch also facilitates the fusion of point cloud features. In contrast, the improved feature extraction from the point cloud branch enhances the fusion for the image branch. This demonstrates that with effective information complementarity between image and point cloud data, both modalities can achieve better perception results.
As illustrated in Fig.5, our approach significantly improves mIoU for small objects like bicyclists, pedestrians and traffic cones. The combination of multiple perspectives and multiple modalities provides more complete information on small objects and improves the segmentation accuracy of small objects.
As illustrated in Fig.6, compared with other light field semantic segmentation methods, our method performs better in the cases of missed segmentation and incorrect segmentation of occluded objects. This is because we fully consider the occlusion problem caused by the occlusion relationship of different modalities and use the ddpm module to perceive the occluded objects.
The point cloud visualization of Mlpfseg on the test set is shown in Fig. 7. The red regions highlight the mispredicted points. The visualization reveals that when the bus is partially occluded, our Mlpfseg method significantly outperforms the baseline. Additionally, for small objects such as pedestrians, our method demonstrates improved accuracy. This performance gain can be attributed to our comprehensive fusion of image and point cloud data, along with the design of the Depth Difference Perception Module (DDPM), which enhances the network’s ability to recognize occluded objects.
IV-A2 Implementation details
Image Semantic Network Implementation: For the input images (1080×1440), we apply several augmentation techniques, including random horizontal flipping (50%), color jittering, JPEG compression noise (quality range [30, 70]), and random cropping (60%-75%). The image segmentation network uses HRNet-W48 as the backbone to extract multi-scale features. We fine-tune a pre-trained HRNet-W48 model on ImageNet, freezing the first three stages. After feature extraction across four stages, the image features $F_{\text{img}}\in\mathbb{R}^{c_{\text{img}}\times h\times w}$ (where $c_{\text{img}}=48$) are fused and passed through a modified FCN head for segmentation. For Mipfseg, we use 3 cameras. Point Cloud Semantic Network Implementation: The point cloud network uses a modified UNet3D architecture with an improved voxel feature extractor. The voxelization occurs within the Cartesian space $x,y,z\in[-50,6,-7]$ to [50, 106, 11], with a resolution of 0.1m and a max of 5 points per voxel. The UNet3D structure applies 8× downsampling and upsampling with a channel scaling factor of 2. After encoding, the point cloud features are $F_{\text{voxel}}^{l}\in\mathbb{R}^{N_{1}\times c_{p}}$, where $c_{p}=48$. Other configurations follow Mseg3D [27]. Training Configuration: Both networks are trained end-to-end using the Adam optimizer, with an initial learning rate of 0.0002 and a weight decay of 0.01. The learning rate follows a one-cycle policy, with a max value of 0.0002 and momentum of [0.95, 0.85]. Training is conducted on a single Nvidia A40 GPU with a batch size of 1 for 24 epochs. A batch size of 1 is also used during inference.
TABLE II: Ablation experiment

### v2

@@ III-B1 Point-Pixel Feature Fusion Module (PFFM)
Considering the projection characteristics of LiDAR point clouds, valid points are mainly concentrated in the central image region, while peripheral areas contain few or no projections. To model this property, we first compute the minimum bounding rectangle $M$ enclosing all projected points on the image plane. The valid projected features within $M$ are denoted as $P_{\text{in}}^{\text{point}}\in\mathbb{R}^{N\times c_{p}}$, while pixels outside $M$ are initialized with zero.
Specifically, let $(x_{\min},y_{\min})$ and $(x_{\max},y_{\max})$ be the minimum and maximum coordinates of the projected points, respectively. The bounding region is defined as $M=[x_{\min},x_{\max}]\times[y_{\min},y_{\max}]$.
@@
$\text{mask}(x,y)=\begin{cases}1&(x,y)\in\{(x_{indices},y_{indices})\}\\ 0&(x,y)\notin\{(x_{indices},y_{indices})\}\end{cases}$ (3)
@@
For each unassigned point $(x,y)$, we find its three nearest valid points $\{(x_{i},y_{i})\}_{i=1}^{3}$, calculate the interpolation weights, where the weights are inversely proportional to the distance between the unassigned point and the valid points $w_{i}=\frac{1}{d_{i}+\epsilon}$, where $d_{i}=\sqrt{(x-x_{i})^{2}+(y-y_{i})^{2}}$. We normalize the weights as $\hat{w}_{i}=\frac{w_{i}}{\sum_{j=1}^{3}w_{j}}$. The interpolated features are then:
$\hat{F}_{point}^{img}=\sum_{i=1}^{3}\hat{w}_{i}\cdot F_{point}^{img}(x_{i},y_{i})\$(4)
@@
$\hat{F}^{img}_{point}(x,y)=\begin{cases}\sum_{i=1}^{3}w_{i}\,F^{img}_{point}(x_{i},y_{i}),&\text{if }\text{mask}(x,y)=1\\ F^{img}_{point}(x,y),&\text{otherwise}\end{cases}$ (5)
where
$w_{i}=\frac{d_{i}+\epsilon}{\sum_{j=1}^{3}d_{j}+\epsilon}.$ (6)
This gives the point cloud projection in the feature map of the image plane within the bounding rectangle $M$. For regions outside the rectangle, we fill them with the corresponding image feature $F_{img}$ to obtain the complete point cloud projection feature map filled point $F_{fill\_point}$ is given by $F_{img}(x,y)$ if $(x,y)\notin M$, and $\hat{F}_{point}^{img}(x,y)$ if $(x,y)\in M$. Since the point cloud and image data features are in different spaces and have different network structures, they are not in the same feature space. To facilitate alignment of the point cloud projection feature map with the image feature space for full fusion, we design an alignment loss function (Pixel2point Loss). This loss function minimizes the difference between the feature spaces of the point cloud and image, enabling better fusion of the two. The alignment loss uses Mean Squared Error (MSE) to measure the difference between the point cloud feature map and the image feature map:
$\mathcal{L}_{align}=\frac{1}{N}\sum_{(x,y)}\left\|F_{fill\_point}(x,y)-F_{img}(x,y)\right\|_{2}^{2}\$(7)
We further refine the fused feature map using a self-attention mechanism to capture both intra- and inter-modal dependencies, yielding the final representation $\hat{F}_{\text{fused}}$. Specifically, Query, Key, and Value are projected from $F_{\text{fused}}$ as:
$Q=F_{\text{fused}}W_{Q},\quad K=F_{\text{fused}}W_{K},\quad V=F_{\text{fused}}W_{V},$ (8)
where $W_{Q}\in\mathbb{R}^{C\times C_{q}}$, $W_{K}\in\mathbb{R}^{C\times C_{k}}$, and $W_{V}\in\mathbb{R}^{C\times C_{v}}$ are learnable matrices. The attention is computed as:
$\text{Attention}(Q,K,V)=\text{Softmax}\left(\frac{Q^{T}K}{\sqrt{C_{k}}}\right)V^{T}.$ (9)
@@
$\hat{F}_{fused}=\text{LayerNorm}\left(\text{Reshape}\left(\text{Attention}(Q,K,V)\right)\right),$ (10)
@@
$L_{fused\_img}=CE(\hat{y}_{fused},y_{gt})\ ,$ (11)
Figure 5: Overview of the proposed Depth Difference Perception Module (DDPM). The module computes the discrepancy between image-derived depth and point-cloud depth to identify occlusion-induced conflicting regions and guides feature fusion accordingly.
@@ III-B2 Depth Difference Perception Module (DDPM)
Occlusion introduces inconsistent observations across modalities, where corresponding regions in images and point clouds may contain conflicting geometric cues. A single modality is often insufficient to resolve such ambiguity. To address this issue, we propose a Depth Difference Perception Module (DDPM), as illustrated in Fig. 5, which explicitly captures the discrepancy between image-estimated depth and LiDAR depth to identify occluded regions.
Depth Discrepancy Modeling. Given light field images $\{\mathcal{L}_{1},\mathcal{L}_{2},\cdots,\mathcal{L}_{n}\}$, we employ ZoeDepth [45] to predict dense depth maps $D_{\text{pred}}$. Meanwhile, LiDAR points are projected onto the image plane according to Eq. (2), yielding sparse depth values $z_{i}$ at pixel locations $(u_{i},v_{i})$. This produces a sparse depth map $D_{\text{sparse}}$, where valid entries correspond to projected LiDAR points.
We only consider valid projection locations $\mathcal{V}$. The depth discrepancy is computed as:
$D_{\text{diff}}(i,j)=\log\left(D_{\text{pred}}(i,j)+\epsilon\right)-\log\left(D_{\text{sparse}}(i,j)+\epsilon\right),$ (12)
where $(i,j)\in\mathcal{V}$ and $\epsilon=10^{-8}$ ensure numerical stability. The logarithmic transformation emphasizes discrepancies at near distances while suppressing noise at far ranges.
Feature Enhancement. To align the depth discrepancy with the feature space of the network, we apply a lightweight two-layer convolution:
$\hat{D}_{\text{diff}}=\text{Conv}(\text{Conv}(D_{\text{diff}})).$ (13)
The refined discrepancy feature $\hat{D}_{\text{diff}}$ is then integrated into the attention mechanism:
$\text{Attention}_{\text{final}}=\text{Attention}(Q,K,V)+\hat{D}_{\text{diff}},$ (14)
which guides the network to focus on geometrically inconsistent regions caused by occlusion.
Loss Functions. For the image branch, the segmentation loss is defined as:
$L_{\text{img}}=CE(\hat{y}_{\text{img}},y_{\text{gt}}).$ (15)
For a single-view input, the total image loss is:
$L_{\text{img\_total}}=L_{\text{img}}+L_{\text{img\_lovasz}}+L_{\text{fused\_img}}+\mathcal{L}_{\text{align}}.$ (16)
For multi-view inputs, the loss becomes:
$L_{\text{img\_total}}=$ $\;L_{\text{img\_center}}+L_{\text{img\_lvcenter}}+L_{\text{img\_lovasz}}+\mathcal{L}_{\text{align}}$ (17)
$+L_{\text{fused\_img}}+\alpha_{1}\sum_{i=1}^{n}L_{\text{img}_{i}}+\alpha_{2}\sum_{i=1}^{n}L_{\text{img\_lvi}},$
where $L_{\text{img\_lvcenter}}$ denotes the Lovász-Softmax loss applied to the center-view prediction, and $L_{\text{img\_lvi}}$ is its multi-view extension for each sub-view.
For the point cloud branch, the losses are defined as:
$L_{\text{point}}=CE(\hat{y}_{\text{point}},y_{\text{point}})+L_{\text{lovasz}}(\hat{y}_{\text{point}},y_{\text{point}}),$ (18)
$L_{\text{voxel}}=CE(\hat{y}_{\text{voxel}},y_{\text{voxel}})+L_{\text{lovasz}}(\hat{y}_{\text{voxel}},y_{\text{voxel}}),$ (19)
$L_{\text{point\_total}}=L_{\text{point}}+L_{\text{voxel}}.$ (20)
The overall training objective is:
$L_{\text{total}}=L_{\text{img\_total}}+L_{\text{point\_total}}.$ (21)
@@ IV Experimental Results
To validate the effectiveness of our proposed the multimodal fusion-based segmentation approach, we conduct experiments on the TrafficScene dataset. All experiments were conducted on a server with an Intel 6330 CPU, 1.0 TB memory, Ubuntu 22.04.5 and CUDA version 12.2. The dataset is split into training, validation, and test sets at a 7:1:2 ratio (3924/594/1116 light field images, 436/66/124 point clouds), using stratified sampling to balance category distributions.
@@
CMNeXt [29] $\times$ $\times$ $\checkmark$ 83.61
@@
Mlpfseg (light field images) $\times$ $\checkmark$ $\checkmark$ 84.97(+3.75) 92.38(+2.38)
Figure 6: mIoU for PSPNet, PSPNet_LGA and Mlpfseg on small objects across all viewing angle
((a))
((b))
Figure 7: comparison of occluded object segmentation in challenging traffic scenes. The top example shows an occluded bicycle, where PSPNet_LGA [10] fails to recognize the target object under heavy occlusion. The bottom example shows an occluded pedestrian, where PSPNet_LGA suffers from incorrect identification, while the proposed Mlpfseg produces robust and accurate segmentation results.
@@ IV-A Semantic segmentation algorithm based on multimodal data fusion
IV-A1 Implementation Details
[…]
IV-A2 Experimental Results
To validate the effectiveness of our dataset and multimodal fusion method, we conducted extensive experiments. First, we assessed the dataset’s validity by applying established image and light-field semantic segmentation methods, evaluating performance using mean intersection over union (mIoU). We conduct extensive evaluations against representative methods across different modalities, including image-based models (e.g., PSPNet [22], DeepLabV3 [23], SegFormer [26]), light-field-based methods (PSPNet_LGA [10], CMNeXt [29]), LiDAR-only methods (SPVCNN [36], MinkowskiNet [35]), and multimodal fusion approaches (PMF [38], 2DPASS [7], MSeg3D [8], DGFusion [46]). All methods are evaluated under identical settings for a fair comparison, and results are reported in Table I.
@@ IV-A2 Experimental Results
The results show that attention-based models consistently outperform convolution-based ones in single-image segmentation. Among them, SegFormer [26] achieves the best performance with 83.15 mIoU. For light-field segmentation, CMNeXt [29] further improves performance by leveraging multi-view sub-aperture information, reaching 83.61 mIoU. For LiDAR-based segmentation, DGFusion [46] achieves the best performance among all compared methods, demonstrating the effectiveness of image-guided fusion. In contrast, projection-based methods such as PMF [38] suffer from 3D structural information loss, resulting in inferior performance compared to point-cloud-only methods such as SPVCNN [36] and MinkowskiNet [35].
@@
Our proposed Mlpfseg consistently outperforms existing state-of-the-art methods in both image and point cloud segmentation, benefiting from image feature interpolation, an attention-based image encoder, and the proposed Depth Difference Perception Module (DDPM) for occlusion modeling. The performance gains can be attributed to three key factors: image feature interpolation alleviates the sparsity caused by point cloud projection, the attention-based image encoder enhances discriminative feature extraction, and the DDPM improves occluded object modeling by explicitly capturing cross-modal depth discrepancies.
As shown in Fig. 6, Mlpfseg achieves significant improvements on small and thin objects such as bicyclists, pedestrians, and traffic cones. The integration of multi-view and multimodal information provides more complete contextual and geometric cues, leading to more accurate segmentation of small objects. Fig. 7 demonstrates that our method effectively handles occluded objects, significantly reducing both missed detections and misclassifications compared with light-field-based methods. This improvement is largely attributed to the proposed DDPM, which enhances the network’s ability to perceive occlusion relationships across modalities. The point cloud results in Fig. 8 show that Mlpfseg produces more accurate predictions under challenging scenarios such as partial vehicle occlusion and small object segmentation. The reduction of mispredicted regions highlighted in red further confirms the effectiveness of the proposed multimodal fusion strategy.
IV-A3 Generalization on Public Datasets
To evaluate generalization, we conduct experiments on two standard benchmarks, SemanticKITTI [47] and nuScenes [48]. SemanticKITTI provides 22 sequences with 19 semantic classes, while nuScenes contains 1000 driving scenes with RGB–LiDAR data and 17 categories.
Following standard protocols, we report point cloud mIoU as the evaluation metric. Since these datasets do not provide light-field inputs, Mlpfseg is evaluated using a single RGB view together with LiDAR for a fair comparison with existing RGB–LiDAR methods. The results are summarized in Table II.
TABLE II: Comparison with state-of-the-art methods on the SemanticKITTI and nuScenes datasets.
Method SemanticKITTI[47] nuScenes[48]
SPVCNN [36] 66.4 77.3
Cylinder3D [5] 64.9 76.1
2DPASS [7] 70.1 80.5
PMF [38] 63.9 76.0
Mseg3D [8] 70.4 81.1
PASeg [49] 70.5 80.5
Baseline 67.9 80.5
Mlpfseg (one view) 71.8 81.6
TABLE III: Ablation experiment